### v1

# Rubric Is All You Need: Enhancing LLM-based Code Evaluation With Question-Specific Rubrics

Abstract.
Since the disruption in LLM technology brought about by the release of GPT-3 and ChatGPT, LLMs have shown remarkable promise in programming-related tasks. While code generation remains a popular field of research, code evaluation using LLMs remains a problem with no conclusive solution. In this paper, we focus on LLM-based code evaluation and attempt to fill in the existing gaps. We propose multi-agentic novel approaches using question-specific rubrics tailored to the problem statement, arguing that these perform better for logical assessment than the existing approaches that use question-agnostic rubrics. To address the lack of suitable evaluation datasets, we introduce two datasets: a Data Structures and Algorithms dataset containing 150 student submissions from a popular Data Structures and Algorithms practice website, and an Object Oriented Programming dataset comprising 80 student submissions from undergraduate computer science courses. In addition to using standard metrics (Spearman Correlation, Cohen’s Kappa), we additionally propose a new metric called as Leniency, which quantifies evaluation strictness relative to expert assessment. Our comprehensive analysis demonstrates that question-specific rubrics significantly enhance logical assessment of code in educational settings, providing better feedback aligned with instructional goals beyond mere syntactic correctness.
Large Language Models, Code Assessment and Grading
*Equal Contribution
journalyear: 2025copyright: acmlicensedconference: 2025 ACM Conference on International Computing Education Research; August 3–6, 2025; Charlottesville, Virginia, United Statesprice: 15.00ccs: Social and professional topics Student assessmentccs: Computing methodologies Natural language processing
1. Introduction
The integration of LLMs into computing education has led to groundbreaking advancements, transforming both students and educators experiences (Hellas et al., 2024; Joshi et al., 2023; Hicke et al., 2023; Raihan et al., 2024). In computing education particularly, LLMs have demonstrated potential in generating instructional content (Sarsa et al., 2022; Phung et al., 2023), providing personalized tutoring and doubt solving(Liffiton et al., 2023; Kotalwar et al., 2025) as well as assisting with code evaluation and grading (Joshi et al., 2024; Alkafaween et al., 2024; Phung et al., 2023). Despite these advancements, there are significant gaps in using LLMs for code evaluation and grading. Most studies focus on generating feedback, helping students identify errors and improve their work (Alkafaween et al., 2024; Pankiewicz and Baker, 2023; Vaithilingam et al., 2022). Existing studies have demonstrated LLMs’ ability to provide meaningful insights into code quality, syntax, and logic (Pankiewicz and Baker, 2023; Anishka et al., 2024). Yet, grading, a critical and labor-intensive responsibility for instructors (Mandernach et al., 2013; Normann et al., 2023), has received comparatively little attention (Kiesler et al., 2023; Jacobs and Jaschke, 2024). The evaluation and grading of student code is a critical component of computing education, as it helps instructors assess students’ understanding of programming concepts and computational logic. This process consumes significant time and energy for instructors (Mandernach et al., 2013; Normann et al., 2023), making it a prime candidate for automation.
Automated grading systems have been in use for a while (Ihantola et al., 2010; Messer et al., 2024), but their effectiveness has often been limited due to their reliance on rigid test cases and simplistic evaluation criteria. Prior work in the domain of automated grading has focused on either introductory programming courses or on short answer evaluation (Raihan et al., 2024; Alkafaween et al., 2024; Yousef et al., 2024). Moreover, we note that existing work on feedback and grading of programming assignments has focused on question-agnostic (QA) rubrics (Phung et al., 2023; Fan et al., 2025), emphasizing generic criteria such as correctness and syntax across diverse problems, but in reality we find that the human instructors actually use question-specific (QS) rubrics.
In this paper, we focus on evaluating the effectiveness of LLMs in grading student code using question-specific rubrics in more advanced computing courses, such as Object-Oriented Programming (OOP) and Data Structures and Algorithms (DSA), where complex problem-solving demands greater precision and context-awareness. We hypothesize that such an approach utilizing question-specific rubrics would yield more accurate evaluations compared to question-agnostic methods. Thus, the primary research question we seek to answer in this paper is:
”How effectively can an LLM-based grader evaluate student code using question-specific rubric compared to question-agonistic rubric? What measurable differences exist in evaluation quality and feedback specificity between these two approaches?”
We propose three novel techniques for code evaluation: (1) Complete Rubric Evaluation (CRE), is a LLM-agent which assesses student submissions against the entire rubric, prioritizing logical correctness and intentionally overlooking syntax errors, to focus primarily on conceptual understanding. A deterministic compiler-equipped agent is used for checking syntactical correctness ; (2) Pointwise Rubric Evaluation (PRE), is similar to CRE, but evaluates submissions by individually checking each criterion within the rubric, providing detailed and granular feedback. It is comparatively more resource-intensive; and (3) Ensembling Method Evaluation (EME), which enhances reliability by aggregating rubric-based evaluations through majority voting and related ensemble mechanisms, ensuring robust feedback and assessment consistency. Additionally, we introduce a new evaluation metric called Leniency, which measures the strictness or leniency of automated evaluations compared to expert human assessments, providing insights into the relative evaluation rigor.
Our empirical findings demonstrate that question-specific rubrics substantially outperform question-agnostic rubrics, leading to improved accuracy, feedback relevance, and alignment with educational objectives.
Conducting this research required a suitable dataset containing student code submissions, model solutions, grading rubrics and feedback. We found that no such dataset was available publicly. To fill this gap, we created our own dataset, drawing from student submissions for OOP and DSA programming exercises. This dataset includes: (1) problem description, (2) student-submitted code, (3) model solutions, (4) grading rubrics, and (5) qualitative feedback, all of which are essential for benchmarking the performance of LLMs in code evaluation tasks. We will be releasing this dataset soon to the public to facilitate further research in this area. The main research contributions of this work are as follows:
  • We present a new dataset that comprises submissions from two important courses in undergraduate computing education: (1) Object-Oriented Programming (OOP), and (2) Data Structures and Algorithms (DSA). (§3)
  • We introduce three novel techniques for code evaluation and grading: (1) Complete Rubric Evaluation (CRE), (2) Pointwise Rubric Evaluation (PRE) and (3) Ensembling Method Evaluation (EME). (§4)
  • We present a new metric that measures the strictness or leniency of an evaluation system based on rubric. The proposed metric, Leniency, is relative and requires baseline data for comparison. (§5)
  • We perform a comprehensive evaluation of the proposed techniques showing that our proposed techniques outperform all other techniques on both the datasets, achieving high correlation with human graders. (§6)
By demonstrating the effectiveness of question-specific rubrics in LLM-based grading systems, we pave the way for more accurate and efficient automated evaluation tools. These tools have the potential to save instructors significant time and effort, allowing them to focus on other aspects of teaching and mentoring. Additionally, students stand to benefit from more detailed and contextually relevant feedback, which can help them identify areas for improvement and deepen their understanding of programming concepts.
2. Related Work
Effective assessment and feedback mechanisms are fundamental to programming education, serving as critical scaffolds that allow for guiding students through the problem-solving process with proper guidance. Prather et al. (Prather et al., 2017) emphasize that effective feedback extending beyond the binary correct/incorrect judgments based on test-cases, generally provided by traditional automated grading systems is crucial for fostering self-regulated learning. Similarly, Hao and Tsikerdekis(Hao and Tsikerdekis, 2019) note that well-designed feedback promotes metacognitive development. The increasing enrollment in programming courses has rendered grading the programming assignments increasingly labor-intensive and time consuming(Liu et al., 2019). To address this challenge, instructors frequently rely on automated grading tools(Keuning et al., 2018).
2.1. Limitations of Traditional Auto-grading Approaches
Conventional autograding systems have primarily relied on predefined test cases to evaluate student submissions. Lobb and Harlow(Lobb and Harlow, 2016) developed Coderunner, an early system that executed student code and compared outputs against expected results derived from instructor-defined test cases. While this approach allowed for basic assessment of functional correctness, it presented several limitations. Keuning et al. (Keuning et al., 2016) identified that these systems often lacked the ability to understand underlying semantic errors or provide nuanced explanations.
A significant limitation of test-case-based approaches is the burden placed on instructors to create comprehensive test suites manually. Keuning et al (Keuning et al., 2016) highlighted the challenges associated with generating test cases that effectively cover edge cases and potential error conditions. Furthermore, as pointed out by Vreda Pieterse (Pieterse, 2013); traditional systems often focus exclusively on functional correctness, neglecting aspects such as code style, efficiency, and adherence to best practices. This results in an incomplete assessment of student programming abilities and fails to promote holistic coding skills development.
Our approach addresses these challenges by employing instructor-defined rubrics that provide specific evaluation criteria, enabling more meaningful feedback beyond binary judgments while maintaining scalability. Though rubrics are also manually intensive, they have great upside in helping instructors evaluate students consistently and objectively (Chowdhury, 2018), while providing students with clear expectations and constructive feedback to identify their strengths and weaknesses (Otero Avila et al., 2019).
2.2. Emergence of LLMs in Programming Assessment
The emergence of Large Language Models (LLMs) marks a significant paradigm shift in autograding approaches. Recent studies by Denny et al. (Koutcheme et al., 2025) have outlined how LLMs can leverage their extensive training data to grasp nuances of code, identify common errors, and explain complex concepts in an accessible manner.
Phung et al. (Phung et al., 2023) investigated the possibility of LLMs mimicking the style and quality of feedback provided by human tutors, benchmarking ChatGPT(OpenAI, 2020) and GPT-4(OpenAI, 2023) against human instructors in programming education contexts. Their results showed that advanced models like GPT-4 can approach human-like assessment capabilities in specific scenarios, though significant gaps remain in others. Azaiz et al. (Azaiz et al., 2024) examined the effectiveness of GPT-4 in generating feedback, indicating its capability to provide insights beyond basic correctness checks.
Beyond simply evaluating code correctness, LLMs demonstrate potential for enhancing feedback quality. Leinonen et al. (Hellas et al., 2023) analyzed techniques for using LLMs to improve programming error messages, making them more informative and actionable for students. Kiesler et al. (Kiesler et al., 2023) explored how LLMs can provide formative feedback on programming assignments, suggesting that these models can offer more detailed and constructive guidance compared to traditional automated systems.
Though the above work highlights the potential of LLMs in the field, many studies also raised questions on the reliability and consistency of LLM-based evaluation (Seo et al., 2025) While the majority of the cited work focuses purely on providing effective feedback, we comprehensively also cover grading assignments which is essential for both instructors and students, especially for a large-scale programming course. Our approach using question-specific rubrics also ensures structured evaluation criteria that combine human expertise with AI capabilities to provide consistent, contextual feedback.
2.3. Major Frameworks and Approaches
Several frameworks have emerged to harness LLMs for programming assessment, and we are focusing on the ones mentioned in Table 1. CodeBERTScore (Zhou et al., 2023) marked a significant milestone in code evaluation by leveraging pre-trained BERT (Devlin et al., 2019) models to encode semantic vectors of reference and generated code, measuring the similarity between these vectors rather than relying on token-matching approaches. This method improved upon earlier techniques like traditional BLEU (Papineni et al., 2002) (which treated code as mere text) and its specialized derivative CodeBLEU (Ren et al., 2020), which had begun addressing code’s unique properties through weighted n-gram matching, AST comparison, and data-flow analysis. Despite CodeBERTScore’s innovations in capturing semantic meaning, it still faces a fundamental limitation: context similarity doesn’t necessarily represent semantic similarity, resulting in suboptimal performance when evaluating functionally identical code implemented with different approaches—a particular challenge in educational environments.
Recent work has explored using LLMs directly for code evaluation without relying on test cases. ICE-Score (Zhuo, 2024) pioneered this approach, making it more applicable in scenarios where test development is impractical. Despite its innovation, this approach demonstrates limited correlation with human judgment and remains susceptible to LLM hallucinations, particularly when evaluating complex code with intricate semantics.
Similarly, CodeJudge (Tong and Zhang, 2024) presents a promising approach leveraging ”slow thinking” to guide LLMs in evaluating code semantics. By decomposing evaluation into step-by-step analysis followed by summarization, and introducing a taxonomy of code inconsistencies with severity levels, it achieves better correlation with semantic correctness compared to existing methods. However, both approaches lack the context of specific programming assignments, limiting their effectiveness for educational assessment.
The techniques discussed above though have application in education are primarily used as benchmarks for code generation by LLMs for task description hence lack context and are unable to generate any meaningful feedback. Beyond these, we focused on rubric-based techniques for code evaluation and found a huge gap in methods exploring the same. For ease of our discussion, we further subdivide rubric-based programming evaluation on the basis of the Rubric specificity into a generic or question-agnostic rubric against a question-specific rubric accounting for the logical requirements from the problem.
While rubrics provide significant benefits for assessment, question-agnostic rubrics face challenges, particularly with automated grading systems for programming assignments. These general rubrics often fail to capture the nuances of specific programming problems, leading to misaligned evaluations, still struggling with grading feedback compared to human tutors (Phung et al., 2023). Their experiments showed that both GPT-4 and ChatGPT frequently misidentify code issues and incorrectly assign points for general correctness and edge cases when using general rubrics, suggesting question-specific rubrics might be more effective for automated assessment systems. More recent developments include domain-specific approaches like CodEV (Tseng et al., 2024), which leverage LLMs, Chain of Thought (CoT), LLM ensembles and a question agnostic rubric to improve score accuracy and consistency while agreement tests help ensure reliable feedback and precise code review comments.
Xie et al (Xie et al., 2024) propose a systematic approach to rethinking automated assessment for short answer with LLMs through their ”Grade Like a Human” system. This multi-agent system divides the grading process into three key stages: (1) developing context-aware grading rubrics that consider both the questions and student answers, (2) providing accurate scoring with customized feedback under these rubrics, and (3) conducting post-grading review to ensure assessment fairness and accuracy. It was designed primarily for short answer grading and hence not directly applicable to the programming assignment evaluation paradigm.
Building on previous methods, our approaches distinctly separate logical reasoning from syntax evaluation, offering practical guidelines for educational contexts where logical understanding outweighs syntactic correctness. Our multi-tiered rubric framework addresses programming education needs by simulating human grading practices, prioritizing conceptual understanding over minor implementation details. Our approaches also differentiate themselves by introducing a systematic rubric granularity framework— from complete to pointwise evaluation —allowing instructors to control assessment strictness based on pedagogical needs.
Table 1. Comparison of Code Assessment Techniques
Technique Grading Feedback
Programming
Focused
Question
Agnostic
Rubric
Question
Specific
Rubric
BLEU(Papineni et al., 2002)* $\times$ $\times$ $\times$ $\times$
CodeBLEU* (Ren et al., 2020) $\times$ $\times$ $\times$
CodeBERTScore*(Zhou et al., 2023)* $\times$ $\times$ $\times$
ICE-Score* (Zhuo, 2024) $\times$ $\times$ $\times$
CodeJudge* (Tong and Zhang, 2024) $\times$ $\times$
Phung et al(Phung et al., 2023) $\times$
CodEv (Tseng et al., 2024) $\times$
Proposed Techniques
*Focused on evaluating LLM-based code generation for task description
3. Dataset
We present two datasets, one based on OOP and second based on DSA. These courses are taken by all Computer Science undergraduates making them relevant to our study. Overall, our dataset contains 230 student submissions (80 from OOP and 150 from DSA). Full Dataset is available on Github  [^1] https://github.com/codeeval18/rubric-code-eval. Below, we provide the high level details about the dataset:
3.1. OOP Dataset Construction
For the OOP dataset, we rely on student submissions as part of a programming exam conducted in a OOP course at University X in fall 2024. The programming exam consisted of one Java programming question which further consisted of seven methods which needed to be implemented by the students. Students were provided with a structured starter template that included class declarations, helper functions, and predefined method scopes for implementation. A main function was also supplied to the students for testing and validating their solutions. The tasks in these methods focussed on applying object-oriented programming principles to handle file I/O, data filtering, and data updates.
We selected 80 student submissions from a total of 350 submissions, all from undergraduate sophomores. The submissions were graded by Teaching Assistants (TAs) and categorized into four score ranges: 0–10, 10–20, 20–30, and 30–35 marks (max score was 35). From each category, 20 solutions were selected at random, totaling to 80 student submissions. The dataset consists of the following components:
  1. (1)
    Detailed Problem Statement: A detailed problem statement provided by the professor, outlining clear, step-by-step instructions for implementing each function in every part of the question.
  2. (2)
    Scaffold Code: The code provides a foundational class structure with templates and clear markers indicating where students should insert their code.
  3. (3)
    Rubric: The rubric outlines each step as a distinct component, assigned either 1 or 2 marks, with no partial grading. A 2-mark step is awarded either 0 or 2 marks. Each of the seven subproblems follows a structured sequence, with steps arranged in a logical order based on their implementation in the codebase.
  4. (4)
    Model Solution: A model solution prepared by the professor, which includes implementations of all steps while strictly adhering to the guidelines in the problem statement and rubric.
  5. (5)
    Student Submissions: Each student submission consists of three parts:
    • Student code: The student’s submitted Java code.
    • Grades: Part-by-part evaluation conducted through a consensus-driven approach by two graders.
    • Feedback: For each of the seven subproblems, a comprehensive feedback was created collaboratively by both graders based on the problem statement and rubric. Both graders are final-year students with extensive experience in programming courses, ensuring a thorough and consistent assessment process.
CricketDataHandler: readPlayersFromFile Method [9 marks]
Write code for reading player data from the input CSV
file and creating
a list of Player objects.
• Step 1: Create an empty list to store player details.[1 mark]
• Step 2: Open the specified file for reading data. [1 mark]
• Step 3: Ignore the first line since it contains the
  column names. [1 mark]
• Step 4: Read each line one by one until reaching the end of the
  file. [1 mark]
• Step 5: Split the line into different pieces of information.
  [1 mark]
• Step 6: Create a new player using this information.[1 mark]
• Step 7: Add the new player to the list. [1 mark]
• Step 8: Close the file after reading all data. [1 mark]
• Step 9: Return the complete list of players. [1 mark]
1public List<player> readPlayersFromFile(String fileName) throwsIOException {
2List<player> players=new List<player>();
3Scanner sc= null;
4Sc = new Scanner(new FileInputStream("InputCricketData.csv"));
5sc.nextLine();
6while(sc.hasNext()){
7 Player p = new Player();
8 String a=sc.nextLine();
9 String s1=a.split(",")[0];
10 String s2=a.split(",")[1];
11 String s3=a.split(",")[2];
12 String s4=a.split(",")[3];
13 String s5=a.split(",")[4];
14 p.setPlayerName(s1);
15 p.setRole(s2);
16 p.setRunsScored(Integer.parseInt(s3));
17 p.setWickets Taken (Integer.parseInt(s4));
18 p.setTeamName(s5);
19 players.add(p);}
20 sc.close();
21 return players;}
Figure 1. Sample Question from OOP dataset and corresponding student submission for the same
As an example, in the above student code, the line List¡Player¿ players = new List¡Player¿(); is incorrect, as List is an interface and cannot be instantiated directly (Step 1). However, all subsequent steps are correct. The final score, determined by human annotators, is 8 out of 9, with individual step-wise marks assigned as 0, 1, 1, 1, 1, 1, 1, 1, 1.
3.2. DSA Dataset Construction
To ensure diversity and comprehensiveness, we selected DSA problems from the Geeks for Geeks (GFG) practice website [7], spanning 9 topics and 3 difficulty levels (easy, medium, and hard). The distribution of problems across topics and difficulty levels is shown in Table 3 and Figure 3.
Topic Number of questions
Arrays 4
Binary Search 3
Bit Magic (Bitwise Operators) 2
Dynamic Programming 2
Graphs 3
Hash 3
Linked Lists 2
Strings 4
Trees 2
Figure 2. Distribution of Problems by Topic
[…]
Figure 3. Distribution of Problems by Difficulty Level
For each problem, the dataset contains:
  1. (1)
    Problem Statement - The problem statements were sourced from the descriptions provided on the website for each question. These include textual description of the problem, input size constraints and example input-output pairs to illustrate expected behavior. A sample problem statment is shown in fig 5
  2. (2)
    Model Solution - One model solution was taken from the editorial given with the problem.
  3. (3)
    Submissions - Six distinct submissions were selected for each problem, representing different categories of outcomes : Correct (3 solutions), Wrong (1 solution), TLE (1 solution) and Compilation error (1 solution).
  4. (4)
    Rubric - The rubric was designed by considering different approaches to solving the question. Marks were assigned to each step based on its relevance and significance across various possible solutions, a sample rubric is shown in fig 6
  5. (5)
    Feedback and Marking - Two human graders collaboratively evaluated six different solutions per problem, following the predefined rubric. They also provided detailed feedback for each implementation step as shown in fig 7. Both graders, as final-year students with extensive programming experience, ensured a thorough and consistent assessment process.
A sample comprising student solution is shown below:
1class Solution {
2 // Function is to check whether two strings are anagram of each other or not.
3 public static boolean areAnagrams (String s1, String s2) {
4 // Your code here
5 if(s1.length() != s2.length() )
6 return false;
7 int charcount[] = new int[256];
8 for(int i = 0; i < s1.length(); i++ ){
9 charCount[s1.charAt(i)]++;
10 charCount[s2.charAt(i)]--;
11 }
12 for(int count: charCount)
13 if( count != 0)
14 return false;
15 return true;
16 }
17}
Figure 4. Sample Student Solution for DSA
Given two strings s1 and s2 consisting of lowercase characters. The task is to check whether two given
strings are an anagram of each other or not. An anagram of a string is another string that contains the same
characters, only the order of characters can be different. For example, "act" and "tac" are an anagram of
each other. Strings s1 and s2 can only contain lowercase alphabets.
Note: You can assume both the strings s1 & s2 are non-empty.
Examples:
Input: s1 = "geeks", s2 = "kseeg"
Output: true
Explanation: Both the string have same characters with same frequency. So, they are anagrams.
Input: s1 = "allergy", s2 = "allergic"
Output: false
Explanation: Characters in both the strings are not same, so they are not anagrams.
Input: s1 = "g", s2 = "g"
Output: true
Explanation: Character in both the strings are same, so they are anagrams.
Constraints: 1$\leq$ s1.size(), s2.size() $\leq$ 105
Figure 5. Sample Problem Statement for DSA
Anagram:
1. Check if two Strings are Anagrams of each other
2. Solution 1:
   1. Initialize a map or dictionary or array to store the character frequencies. Initially, the frequency
      for each character should be zero.[1 mark]
   2. For each character in the first string, the corresponding frequency is incremented by 1. [1 mark]
   3. For each character in the second string, decrement its corresponding frequency by 1. [1 mark]
   4. Iterate through the entire map or dictionary or array. If any frequency is non-zero, then return false. Else return
    true. [1 mark]
3. Solution 2:
   1. Initialize a map or dictionary or array to store the character frequencies. Initially, the frequency
      for each character should be zero. Keep separate data structures for each of the strings. [1 mark]
   2. For each character in the first string, the corresponding frequency in its data structure is
      incremented by 1. [1 mark]
   3. For each character in the second string, the corresponding frequency in its data structure is
      incremented by 1. [1 mark]
   4. Iterate through both the data structures. If the frequency for any unique character in both the
      data structures does not match, return false. Else return true. [1 mark]
4. Solution 3:
   1. Sort both the strings. [2 marks]
   2. Return true if both the sorted strings are exactly same. Else, return false. [2 marks]
Figure 6. Sample Rubric for DSA
"Selected rubric" Solution 1
"Feedback with marks"
1. Initialize a map or dictionary or array to store the character frequencies. Initially, the frequency
    for each character should be zero. [1 mark]
      -The student initializes an array ‘charCount‘ to store character frequencies, which is correct.[1 mark]
2. For each character in the first string, the corresponding frequency is incremented by 1.[1 mark]
      -The student correctly increments the frequency for each character in the first string ‘s1‘ using
       ‘charCount[s1.charAt(i)]+‘. [1 mark]
3. For each character in the second string, decrement its corresponding frequency by 1. [1 mark]
      -The student correctly decrements the frequency for each character in the second string ’s2’ using
       ’charCount[s2.charAt(i)]--’. [1 mark]
4. Iterate through the entire map or dictionary or array. If any frequency is non-zero, then return false.
    Else return true. [1 mark]
      -The student iterates through the ‘charCount array and checks if any frequency is non-zero, returning
       false if so, and true otherwise. [1 mark]
"Marks Scored out of total marks"
4 out of 4 marks
Figure 7. Sample Grade and Feedback for DSA
4. Proposed Techniques
4.1. Motivation
The proposed techniques discussed below were designed in order to evaluate submissions using question-specific rubrics. As discussed earlier, question-specific rubrics were introduced in order to emulate college or university-level grading of student code. Akin to college grading, we break down the problem statement into steps that the student is expected to implement. Marks are awarded with respect to the student’s approach for each specific step. The marking is binary, so a student is awarded either full or zero marks for a step. Our datasets attempt to emulate a similar level of leniency and specificity with grading of the ground truth values.
4.2. Complete Rubric Evaluation (CRE)
Complete Rubric Evaluation (CRE) is a system of rubric-based evaluation developed as part of our research into LLMs and their capabilities in code evaluation. As shown in Figure 8, the CRE LLM grader agent takes as input the complete problem description followed by a complete rubric (a multi-tiered point-by-point marking scheme). The rubric points act as anchors around which the student code is evaluated. Finally, the grader agent takes in the entire student code file, including all methods and classes implemented by the student.
The grader returns a single JSON dictionary without additional text. The JSON dictionary is a nested rubric, where the primary keys are method names. The values may either be the marks assigned to the method or another dictionary containing finer evaluation points for the method. Since LLMs are unreliable in arithmetic operations, the final calculation of marks is performed by a recursive function outside the LLM loop.
The LLM in CRE grader is instructed to ignore syntax errors in the student code, treating them as correct. CRE aims to infer the logical intention behind student code and mark it accordingly. This simulates a university-like evaluation environment, where logical reasoning is prioritized over minor syntax errors. This approach addresses gaps in traditional evaluation methods (e.g., test-case-based evaluation), which assign zero marks to submissions with minor errors.
While the LLM performs only logical evaluation, syntax evaluation is conducted separately using a deterministic compiler-equipped agent. The student code is executed by the agent using a compiler via a system call in a Python script. The compiler returns a syntax assessment, and the agent uses a penalty-based system to assign syntax marks accordingly. For instance, with a maximum of 5 syntax marks and a penalty of 0.5, a code snippet with 5 syntax errors is awarded 2.5 marks. The final student marks are the sum of logical and syntactical scores. Prompts used for instructing LLMs are shared in Appendix A
Figure 8. Complete Rubric Evaluation (CRE)
4.3. Pointwise Rubric Evaluation (PRE)
Pointwise Rubric Evaluation (PRE) is a modified version of CRE. Instead of evaluating the entire rubric at once, the PRE LLM grader agent assesses the student code based on a single rubric point at a time. As shown in Figure 9, PRE Grader takes as input the problem statement, student code, and a single rubric point for evaluation. It returns a JSON dictionary containing the rubric point as the key and the assigned LLM marks as the value. These results are stored in a JSON file for later computations.
PRE is resource-intensive due to multiple API calls per student solution. An LLM call is executed in order to evaluate each point in the rubric, consuming tokens and time.
Figure 9. Pointwise Rubric Evaluation (PRE)
4.4. Ensembling Method Evaluation (EME)
Inspired by CodEv (Tseng et al., 2024), Ensembling Method Evaluation (EME) leverages large language models (LLMs), such as GPT-4o, Claude 3.7 sonnet(cla, [n. d.]) and GPT-4o mini, to validate ensemble-generated results using a structured evaluation framework. The method employs a sampling and voting-based approach, primarily relying on the majority voting method to determine the final ensemble output. In cases where no clear majority emerges, the rounded mean method is used as an alternative to aggregate the scores effectively. Additionally, EME incorporates a feedback selection mechanism, where the system identifies the most representative feedback by selecting the evaluation feedback whose total score is closest to the final ensemble score. As shown in Figure 10, this technique takes the question, the student solution, the reference solution and the rubric. In case of DSA dataset we also add a approach identification prompt which uses the GPT 4o model, to identify the which approach the student has used from the rubric. This prompt provides us with a confidence interval for the approach identified by the model.
Figure 10. Ensembling Method Evaluation (EME)
5. Metrics and Evaluation
5.1. LLM selection
In order to maintain a balance between LLM accuracy and token costs, we used OpenAI’s GPT-4o-mini model to conduct our evaluations. The model has been shown to be remarkably accurate and consistent while maintaining significantly lower API costs versus OpenAI’s other models, including GPT-4o, o1 or o1-mini. We also used Anthropic AI’s Claude 3.7 Sonnet model with Extended Thinking deactivated.
5.2. Data preprocessing
5.2.1. Scaling
With varying scales and degrees of rubrics, the scale at which grades are awarded is different for every technique. Here, scale refers to the maximum marks or score that can be awarded for the evaluation technique (e.g. 35 for the OOP question-specific rubric, or 4 for the question-agnostic rubric). For the purposes of understanding, we shall refer to the Ground Truth values as the Base Dataset (B) and the LLM evaluations as the Experiment Dataset (E). $B[i]$ and $E[i]$ denote the base and experimental marks obtained by the $i^{th}$ student, where $0\leq i<N$. To assess the correspondence of an Experiment Dataset E that lies on a different scale versus B, we scale E linearly as
$E_{s}[i]={E[i]}\times\frac{R}{R_{E}},\forall 0\leq i<N$
where $R_{E}$ (e.g. 4 for the question-agnostic rubric) is the scale of the experimental rubric, $R$ is the scale of the question-specific rubric (also the scale of ground-truth values, e.g. 35 for the OOP dataset), $N$ is the number of data points in the dataset and $E_{s}$ is the scaled experimental dataset. For techniques evaluated method-by-method on the OOP dataset using question-agnostic rubrics, we scaled $E$ on the method level as
$E_{s}[i]=\sum_{m=1}^{M}{(E[i][m]\times\frac{R[m]}{R_{E}})},\forall 0\leq i<N$
where $m$ is the method being evaluated, $M$ is the number of methods present in the problem statement and $R[m]$ are the maximum number of marks that can be awarded for the method in the question-specific rubric. The total marks for a student would therefore be the sum of marks awarded for each method, scaled up to match the base rubric.
5.2.2. Binning
Before we obtain Cohen-Kappa scores on continuous datasets, we first classify the data into 5 bins. The bins are decided based on the ranks of the data points in the sorted datasets. For instance, the dataset $[3,2,4,3,1,9,6,7]$ when classified into 3 bins is graded as $[0,0,1,1,0,2,1,2]$, where 0, 1 and 2 are the three possible grades, assuming uniform grading.
5.3. Evaluation Metrics
In order to assess the performance of question-specific rubrics versus question-agnostic rubrics on the proposed datasets, we shall use the following metrics:
  • Pearson Correlation Coefficient (PCC, $r$) (sci, 2025): PCC measures linearity between two datasets. A value of magnitude 1 indicates perfect linearity between the sets. Positive values indicate an increasing nature in $E$ as $B$ increases, while negative values indicate a decreasing nature.
  • Spearman Rank Correlation Coefficient ($r_{s}$) (sci, 2025): Spearman Correlation measures the degree of monotonicity between the two sets. Similar to the Pearson coefficient, an absolute value of 1 indicates perfect monotonicity, while positive and negative values denote increasing and decreasing trends in $E$ while $B$ increases.
  • Kendall-Tau Correlation Coefficient ($\tau_{b}$) (sci, 2025): Also called the Kendall Rank Correlation Coefficient, it measures ordinal association between two quantities. It is a measure of rank correlation, i.e., it denotes the similarity of the orderings of data points when ranked. Values of positive or negative 1 indicate perfect association, while the sign of the coefficient denotes an increasing or decreasing trend.
  • Leniency (Mean Normalized Error, $l_{n}$): The leniency between two datasets is defined as
    $l_{n}=\frac{\sum_{i=0}^{N-1}\hat{E}[i]-\hat{B}[i]}{N}$
    where $\hat{E}[i]$ and $\hat{B}[i]$ are the experimental and base scores normalized to 1. Leniency is a measure of how strict the evaluation system is for a rubric. Leniency is relative, and requires base data against which the metric is computed. A perfectly lenient grader would assign R marks where every base data point is 0, achieving a leniency score of 1. A perfectly strict grader would assign 0 marks where every base data point is R, achieving a leniency score of -1. If we consider a target average deviation of 10%, the leniency to aim for would be within $\pm$0.1.
  • Intraclass Correlation Coefficient (ICC) (pin, 2024): The ICC is used to assess the consistency of measurements made by multiple observers that measure the same quantity. We measure the ICC of the llm-based evaluation techniques graders against the base dataset $B$ in order to estimate how well the experimental scores match the truth. We use the following 3 ICC scores
    • One-way random effects (ICC1): each subject is measured by a different set of k randomly selected raters;
    • Two-way random (ICC2): k raters are randomly selected, then, each subject is measured by the same set of k raters;
    • Two-way mixed (ICC3): k fixed raters are defined. Each subject is measured by the k raters.
  • Cohen-Kappa w/ Binning (skl, 2024) ($\kappa_{B}$): As described in the previous section, $\kappa_{B}$ is the Cohen-Kappa score obtained after binning the data points into a predetermined number of bins (5 in our case). This is done because Cohen-Kappa measures correspondence between graded scores. Binning allows us to emulate relative grading to an extent.
6. Results
6.1. Techniques Used
  • CodeBERTScore (Zhou et al., 2023): CodeBERTScore is an Automatic Evaluation Metric for Code evaluation, based on BERTScore. We evaluate CodeBERTScore on the OOP dataset and present correlation metrics. Leniency for CodeBERTScore is not evaluated, since the data points lie on a different scale versus the ground truth score.
  • CodeJudge (Tong and Zhang, 2024): CodeJudge is an LLM-based evaluation technique . We implement the binary evaluation module in CodeJudge and evaluate binary scoring for each method. Once the score is obtained, we evaluate complete or zero marks for that method based on the marks in our rubric. CodeJudge achieves comparable results to those presented in the original paper.
  • Ensembling Method Evaluation (QA/QS): We use LLM ensembling to achieve results with a degree of consensus. Similar to CodeJudge, we evaluate each method in the original rubric, but instead of binary marking, we grade based on a 5 point rubric, awarding scores from 0-4. These marks are then scaled to 35 before evaluations.
  • CRE/PRE: Complete and Pointwise Rubric Evaluation are question-specific rubric evaluation techniques discussed earlier in the paper. CRE and PRE are both evaluated on the original 35-mark rubric. For evaluation metrics on OOP, we consider only logical marks, as human grading of the OOP dataset followed a similar logical correctness-based approach.
  • Five Point Marking (FPM) (Phung et al., 2023): Five Point Marking is a question-agnostic LLM-based rubric marking technique. The student code is evaluated based on the following points:
    • Program Format (10 Marks)
    • Time Complexity (15 Marks)
    • Space Complexity (15 Marks)
    • Correctness General (30 Marks)
    • Correctness Edge Cases (30 Marks)
For a total score of 100. The scores are scaled down to 35 for comparison with ground truth values.
6.2. Findings
Tables 2 and 3 provide an overview of the performance of all implemented techniques on the OOP and DSA datasets. Table 2 provides an overview of the evaluation performed on the OOP dataset, with results from EME, CRE and PRE, while Table 3 provides an overview of the results obtained for the DSA dataset. We evaluate the question-agnostic versus question-specific rubrics on EME.
Method $r$ $r_{s}$ $\tau_{b}$ $l_{n}$ ICC1 ICC2 ICC3 $\kappa_{B}$
No LLM (Similarity Based)
CodeBERTScore 0.354 0.482 0.343 - - - - 0.241
No Rubric
CodeJudge 0.717 0.745 0.597 -0.233 0.479 0.541 0.712 0.433
Question-Agnostic Rubric
EME (QA) 0.904 0.909 0.775 -0.071 0.881 0.882 0.904 0.512
FPM 0.844 0.834 0.670 -0.121 0.778 0.784 0.834 0.346
Question-Specific Rubric
EME (QS) 0.900 0.902 0.769 -0.067 0.878 0.880 0.900 0.545
CRE (OpenAI) 0.912 0.906 0.773 0.082 0.880 0.882 0.910 0.598
CRE (Claude) 0.840 0.841 0.696 -0.008 0.841 0.841 0.840 0.572
PRE 0.742 0.795 0.637 -0.329 0.201 0.378 0.679 0.302
Table 2. Results obtained by various techniques on the OOP dataset
Method $r$ $r_{s}$ $\tau_{b}$ $l_{n}$ ICC1 ICC2 ICC3 $\kappa_{B}$
No LLM
CodeBERTScore 0.126 0.058 0.039 - - - - 0.010
No Rubric
CodeJudge 0.423 0.427 0.389 -0.315 0.176 0.272 0.353 0.406
Question-Agnostic Rubric
EME (QA) 0.562 0.510 0.445 -0.098 0.509 0.525 0.560 0.156
FPM 0.470 0.381 0.320 -0.054 0.380 0.388 0.398 0.072
Question-Specific Rubric
EME (QS) 0.825 0.763 0.675 0.0049 0.821 0.820 0.819 0.646
Table 3. Results for Question-Agnostic and Question-Specific rubrics on the DSA dataset
Within the scope of this paper, we have discussed building rubrics, and the differences between question-specific and question-agnostic rubrics. We have also introduced evaluation techniques for question-specific rubrics, and compared the performance of LLM grading with question-specific rubrics versus question-agnostic rubrics.
While PRE is instructed to evaluate only logical correctness, it is observed to be more strict and detail-oriented than CRE. PRE consistently awards lower marks compared to CRE, leading to lower correlation metrics, as human grading of the OOP dataset was relatively lenient. The primary reason for this is that PRE, when given a single rubric point, applies stricter criteria. Since the LLM evaluates only one aspect at a time, it tends to assign zero marks unless the code fully meets the given rubric criteria. PRE is suitable for use cases requiring stringent evaluation and strict rubric adherence.
For EME, we observe the following:
  • Model Parameter Size: The performance of EME improves as the parameter size of the underlying language model increases. Larger models tend to demonstrate enhanced reasoning and evaluation capabilities, resulting in more accurate scoring and feedback generation. Conversely, models with relatively lower parameter counts exhibit greater inconsistencies, leading to lower correlation.
  • Ensemble Size: The effectiveness of the method varies with ensemble size. Initially, a significant increase in correlation coefficients is observed as ensemble size decreases. This trend continues until the ensemble reaches an optimal range—typically around three to four models when employing high-parameter LLMs, such as GPT-4 and Claude 3. Beyond this threshold, performance gains tend to plateau, suggesting diminishing returns with further increase in ensemble size.
We observe clearly that question-specific rubric based evaluation techniques outperform all other techniques with respect to the metrics we have used. EME (QS) outperforms EME (QA) on DSA datasets, achieving high correlation, and performing the best out of all techniques in terms of correlation coefficients, leniency, ICC and $\kappa_{B}$. 3
CRE scores consistently high for all correlation scores, and takes the lead with Pearson Correlation as well as with ICC3 and $\kappa_{B}$. 2 PRE has a leniency of -0.329 against ground truth values. This means that on average, PRE scores $35\times 0.329=11.515$ marks lower than the ground truth. In comparison, CRE has a leniency of 0.081, which in magnitude is considerably lower than that of PRE. This corresponds with our hypothesis that LLMs given a single rubric point will be considerably stricter with the evaluation of the student code.
6.3. Variations in LLM Grading Performance
6.3.1. Dataset
: Our techniques consistently yield better performance on the OOP dataset, achieving higher scores across various evaluation metrics. We observe that QA and QS techniques perform relatively well with the OOP dataset. This can be attributed to the nature of the dataset, which consists of questions that are relatively straightforward to implement, with minimal variation in algorithmic approaches or implementation strategies across different solutions. These questions primarily pose an implementation challenge rather than requiring deep logical reasoning. As a result, advanced language models like GPT-4 and Claude 3.7 perform well even without a predefined rubric. In contrast, their performance declines on the DSA dataset, which demands a deeper understanding of algorithms and their underlying logic. The results indicate that while LLMs effectively evaluate solutions in the OOP dataset, they struggle with the greater algorithmic diversity and complexity present in the DSA dataset, especially when a question-specific rubric is absent.
6.3.2. Presence of LLM Grader
: We observe that LLM-based techniques outperform CodeBERTScore significantly and impressively. CodeBERTScore obtains weak scores across all evaluation metrics. LLMs are vastly more versatile when it comes to understanding context, as well as the varying nature of student approaches to one particular problem.
6.3.3. Presence of Rubric
: We observe that both question-agnostic and question-specific rubric techniques outperform no-rubric techniques. While CodeJudge achieves respectable correlation and $\kappa_{B}$ scores, providing a rubric gives the LLM grader an anchor around which to evaluate or grade the code, thereby increasing performance significantly when a rubric is provided.
6.3.4. Specificity of Rubric
: We observe that question-specific rubric techniques outperform question-agnostic rubric techniques in the DSA dataset, while both question-specific and question-agnostic techniques score high within the OOP dataset. We provided the LLM grader the exact rubric that human graders were given, and observe that the LLM grader is able to follow the marking scheme with a satisfying degree of correctness, as evidenced by the consistently high correlation scores, and the observed leniency scores within the range $0.0_{\pm{0.1}}$.
NOTE: We exclude PRE from this comparison for the reasons described next.
6.3.5. Manner of Feeding
: We observe that the manner in which a rubric is presented to a large language model (LLM) can significantly influence the grading outcomes. In the EME framework, we provide a model solution as part of the prompt. This inclusion likely accounts for the comparable performance of EME(QA) and EME(QS), as object-oriented programming (OOP) student implementations tend to closely align with the provided model solution. Conversely, in the case of data structures and algorithms (DSA) evaluation, the model solution has limited utility, as student implementations exhibit substantial variation. Within EME (QS) and CRE, we feed the LLM the entire rubric at once, after which the LLM grader performs grading in a single LLM call. Within PRE, each rubric evaluation point is fed one-by-one. When fed evaluation points one by one, the LLM is remarkably strict in grading, as evidenced by the $l_{n}$ score of -0.329. We therefore make the claim that to emulate grading on a human-like scale, the LLM grader should be fed the entire rubric as an evaluation context.
7. Limitations and Future Work
Although this study provides valuable insights into rubric-based code evaluation, certain limitations must be acknowledged which also advocate avenues for future work. Firstly, we didn’t explore variability among various LLMs and mainly worked with GPT-4o. Different training methodologies may influence effectiveness and accuracy of code evaluation and a future research could compare multiple models. We also didn’t explore the differences between open-sources and closed-source LLMs. Our analysis is limited to programming questions in Java and do not test other widely-used languages like Python, C++. We focussed on 2 intermediary courses which had single file code assignments. A future research could delve into advanced courses which require multiple files handling. The rubrics which can be used for evaluation, can be of various granularities ranging from low to medium to fine. A detailed study on their effectiveness and comparison is another future avenue for research.
8. Conclusion
In this paper, we explored how effectively can an LLM-based grader evaluate student code using question-specific rubric compared to question-agonistic rubric. To facilitate our evaluation, we introduced two novel datasets focused on DSA and OOP, encompassing solutions of varying correctness levels. Using these datasets, we assessed our proposed techniques against existing approaches using exisitng and a new evaluation metric, Leniency. Our findings highlight the limitations of existing evaluation techniques and demonstrate the potential benefits of question-specific rubrics in enhancing logical assessment of code. This work not only provides new insights into LLM-based code evaluation but also lays the groundwork for future research into refining evaluation methods and expanding dataset availability to improve automated code assessment.
9. Acknowledgments
The authors wish to acknowledge the use of ChatGPT/Claude in the writing of this paper. This tool was used to improve the structure, presentation and grammar of the paper. The paper remains an accurate representation of the authors’ underlying work and novel intellectual contributions.
Appendix A Prompts for Techniques
A.1. CRE
1You are an expert code evaluator, evaluating code submissions for a Java based Object Oriented Programming test at a university level.
2You will be provided with the question and a rubric that describes the criteria for evaluation, with a marking scheme.
3The question is a code sample that the examiner provides, containing a template wherein the student is required to write the code as well as comments and instructions from the examiners end.
4Following this you will be provided with the code submission, along with the response from the Java compiler that runs this code.
5Note that the code may be formatted liberally, the specific positioning of the code within the methods are not important.
6Code may be present either before or after the comments prepared by the instructor.
7You are to evaluate the code based only on logical correctness. You are to ignore any syntax errors that the compiler may have thrown.
8Any syntax errors that you encounter can be treated as correct syntax, and you are to infer the students logical flow and intention from the code.
9You are to return your response as a JSON dictionary containing a detailed, nested evaluation of the students marks for each line in the rubric.
10The JSON dictionary should also contain feedback for each point in the rubric.
11For each line in the rubric, you are to provide the line as the key and a nested dictionary containing marks awarded and feedback.
12The following is a sample return Format:
13{
14"1000": {
15 "Method1": {
16 "Point1": {
17 "Marks": 3,
18 "Feedback": {LLM obtained feedback}
19 },
20 "Point2":{...},
21 ...
22 },
23 "Method2: {...},
24 ...
25}
26}
27DO NOT RETURN ANY ADDITIONAL TEXT ASIDE FROM THE JSON DICTIONARY.
28Question: {}
29Rubric: {}
30Code Submission: {}
31Compiler Response: {}
A.2. PRE
1You are an expert code evaluator, evaluating code submissions for a Java based Object Oriented Programming test at a university level.
2You will be provided with the question, the code snippet, and the point of evaluation for the code. You will also be given the compiler response for the code.
3You will also be given the rubric point that the student is graded on. You are to evaluate based on that particular point only.
4The question is a code sample that the examiner provides, containing a template wherein the student is required to write the code as well as comments and instructions from the examiners end.
5Following this you will be provided with the code submission, along with the response from the Java compiler that runs this code.
6Note that the code may be formatted liberally, the specific positioning of the code within the methods are not important.
7Code may be present either before or after the comments prepared by the instructor.
8You are to evaluate the code based only on logical correctness. You are to ignore any syntax errors that the compiler may throw.
9Any syntax errors that you encounter can be treated as correct syntax, and you are to infer the students logical flow and intention from the code.
10You are to return only a dictionary containing the your decision and your feedback, with the keys "DECISION" and "FEEDBACK"
11For your decision, return YES if the student has correctly implemented the logic for the given rubric point, and NO if they have not.
12Since there is no partial marking and were only considering logical correctness, be liberal with the quality of the code and the marking.
13The following is a sample return Format:
14{
15 "DECISION": "YES",
16 "FEEDBACK": {LLM obtained Feedback},
17}
18DO NOT RETURN ANY ADDITIONAL TEXT ASIDE FROM THE DICTONARY.
19Question: {}
20Student Solution: {}
21Point to be evaluated: {}
22Compiler Response: {}
A.3. EME
1## Approach Identification Prompt
2
3You are analyzing a students code submission for a DSA problem.
4Based on the rubric, identify which approach the student is using.
5
6Rubric:
7‘‘‘
8{rubric_content}
9‘‘‘
10
11Student Code:
12‘‘‘
13{code}
14‘‘‘
15
16IMPORTANT INSTRUCTIONS:
171. The rubric contains multiple solution approaches (e.g., "Solution 1", "Solution 2", "Solution 3").
182. Each approach has specific criteria and point allocations.
193. Carefully analyze the students code to determine which approach they are using.
204. Look for key patterns, variable names, and algorithm structures that match one of the approaches in the rubric.
215. The approach name should be EXACTLY as it appears in the rubric (e.g., "Brute Force", "Dynamic Programming", "Kadanes Algorithm").
22
23Respond ONLY with a JSON object in this exact format:
24{
25 "identified_approach": "Exact approach name from rubric (e.g., Solution 1 (Brute Force)’, Solution 3 (Kadanes Algorithm)’)",
26 "confidence": 0.95, // A number between 0 and 1 indicating confidence in the identification
27 "reasoning": "Brief explanation of why you identified this approach, citing specific code patterns that match the rubric criteria"
28}
29
30## System Message for Approach Identification
31You are a code analyzer that ONLY responds with valid JSON. No other text or explanation. You must identify the exact approach from the rubric.
32
33## Code Evaluation Prompt
34
35You are evaluating a students code submission for a DSA problem.
36Provide your evaluation in VALID JSON format only.
37
38Problem:
39‘‘‘
40{question}
41‘‘‘
42
43Rubric:
44‘‘‘
45{rubric}
46‘‘‘
47
48Reference Solution:
49‘‘‘
50{solution}
51‘‘‘
52
53Student Code:
54‘‘‘
55{code}
56‘‘‘
57
58The student appears to be using the "{identified_approach}" approach.
59
60IMPORTANT INSTRUCTIONS:
611. Evaluate the submission according to the EXACT criteria in the rubric for this approach.
622. For each criterion in the rubric for this approach, assign appropriate points.
633. The criterion descriptions should match EXACTLY whats in the rubric.
644. The max_score for each criterion should match the points specified in the rubric.
655. Your feedback should directly address how well the students code meets each specific criterion.
666. Do not create new criteria that arent in the rubric.
67
68Respond ONLY with a JSON object in this exact format:
69{
70 "criteria_scores": [
71 {"criterion": "exact criterion from rubric", "score": awarded_points, "max_score": points_specified_in_rubric, "feedback": "specific feedback for this criterion"}
72 ],
73 "total_score": total_awarded_points,
74 "max_possible_score": total_maximum_points,
75 "overall_feedback": "overall feedback here",
76 "approach_correctness": 0.95, // How confident you are that the approach identification is correct (0-1)
77 "code_correctness": 0.9, // How likely the code is to work correctly (0-1)
78 "efficiency_rating": 0.8, // How efficient the solution is relative to optimal (0-1)
79 "readability_rating": 0.7 // How readable and well-structured the code is (0-1)
80}
81
82## System Message for Code Evaluation
83You are a code evaluator that ONLY responds with valid JSON. No other text or explanation. You must follow the rubric exactly when evaluating code.